# What is and is not inside a Cayley graph?

### v1

@@ 2 No-lonely-color colorings of digraphs
For a group $\Gamma$ and subset $S$ the Cayley digraph $\overrightarrow{\textrm{Cay}}(\Gamma,S)$ has vertex set $\Gamma$, where two elements $a,b\in\Gamma$ form an arc $(a,b)$ if $a^{-1}b\in S$. Note that antiparallell arcs, i.e., of the form $(a,b)$ and $(b,a)$, may occur. We say that $D$ is a minimal Cayley digraph if there is an inclusion-minimal generating set if sone group $\Gamma$ such that $D\cong\overrightarrow{\textrm{Cay}}(\Gamma,S)$. Cleary,
@@
    for every $i\in S$ the subdigraph $f^{-1}(i)$ has maximum outdegree and maximum indegree at most $1$
@@
    all cycles in $f^{-1}(i)$ are directed (by Definition 2.1.1.) and monchromatic (by Definition 2.1.3.).
@@
If $G$ is subgraph of a minimal Cayley graph, then by Observation 2.1 $G$ is the underlying graph of $D$ a subdigraph of a minimal Cayley digraph. Now, we can apply Lemma 2.3 to $D$. By Observatiobn 2.2.2 we get:
@@ 3 An algorithm
[…]
@@
6if all edges are colored and $f\_valid$ then
@@ 25       $f(e)\leftarrow c$
For the following assume we are given $G=(V,E)$ together with a no-lonely-color coloring $f:E\to S$. We devise an algorithm to construct an orientation $D$ of $G$ color by color to obtain a candidate orientation. It is based on the following easy observations. For this, we note that if $H$ is connected and non-trivial, i.e., with more than one vertex of maximum degree $\leq 2$. Then $H$ has exactly two orientations satisfying Definition 2.1.1, denoted by $H^{+}$ and $H^{-}$. By Observation 2.2 in each color, one of these orientations can be arbitrarily fixed. Further, note that by Obervation 2.4, only if some color $f^{-1}(i)$ is a matching, then its edges can potentially be completely bi-oriented. We initialize $D=G$ unoriented and $c=1$.
1 Input: An integer $c$ indicating the next color class to orient.
@@ 12             Bi-orient all edges in $f^{-1}(c)$ in $D$
1 Input:
[…]
@@ 7        $C=(a^{\sigma(1)}_{1},\ldots,a^{\sigma(\ell)}_{\ell})$
8       where each $a_{i}$ is an arc and $\sigma(i)\in\{+1,-1\}$ indicates direction. Let $\mathbf{c}=(f(a_{1}),\ldots,f(a_{\ell}))$ be the color sequence.
9       Let $\boldsymbol{\sigma}=(\sigma(1),\ldots,\sigma(\ell))$ be the direction sequence. for every vertex $v$ in $V$ do
@@
4 Two cubic graphs of grith 5
There are two cubic graphs of grith 5. Our implementation [24] applied to the triplex graph, see the left of Figure 1, shows that it is not the subgraph of a minimal Cayely graph. Indeed, Algorithm 1 produces all 216 faithful no-lonely color colorings associated to this graph, in 11 seconds. After reducing modulo automorphism, we are only left with 15 colorings. Algorithm 2 and Algorithm 3 finish without finding any orientation to any of these colorings in about 45 seconds.
@@ 4 Two cubic graphs of grith 5
The second cubic graphs of grith 5 is the twinplex graph depicted on the right of Figure 1. In this case, our procedure finds an orientation and a coloring that passes our algorithm, as depicted in the figure. Hence, we cannot conclude whether the graph is a subgraph of a minimal Cayley graph or not.
@@ 5 Generalized Petersen Graphs
[…]
@@
$<a,b\mid a^{r}=b^{n}=e,aba^{-1}=b^{k}>.$
We show that $G(n,k)$ is isomorphic to an induced subgraph of $\textrm{Cay}(C_{n}\rtimes C_{r},\{a,b\})$. Indeed, for $10\leq i\leq n-$ 1 map every outer vertex $u_{i}$ to $b^{i}$ and every inner vertex $v_{i}$ to $b^{i}a$. See the left side of Figure 2 for an example. Clearly, by minimality of the generating set, the outer vertices induce a cycle of length $n$ in $\textrm{Cay}(C_{n}\rtimes C_{r},\{a,b\})$. Further, there is an edge between $u_{i}$ and $v_{j}$ if and only if $b^{-i}b^{j}a\in\{a,b\}$ or $a^{-1}b^{-j}b^{i}\in\{a,b\}$. This is the case if and only if $i=j$. Now, consider inner vertices $v^{i}v^{j}$ these have an edge if $a^{-1}b^{-i}b^{j}a\in\{a,b\}$ or $a^{-1}b^{-j}b^{i}a\in\{a,b\}$. Using $aba^{-1}=b^{k}$ one gets that this is the case if and only if $i-j=k\mod n$, which finishes the proof. ∎

### v2

@@ 2 No-lonely-color colorings of digraphs
For a group $\Gamma$ and subset $S$ the Cayley digraph $\overrightarrow{\textrm{Cay}}(\Gamma,S)$ has vertex set $\Gamma$, where two elements $a,b\in\Gamma$ form an arc $(a,b)$ if $a^{-1}b\in S$. Note that antiparallel arcs, i.e., of the form $(a,b)$ and $(b,a)$, may occur. We say that $D$ is a minimal Cayley digraph if there is an inclusion-minimal generating set if sone group $\Gamma$ such that $D\cong\overrightarrow{\textrm{Cay}}(\Gamma,S)$. Clearly,
@@
    for every $i\in S$ the subdigraph $f^{-1}(i)$ has maximum outdegree and maximum indegree at most 11
@@
    all cycles in $f^{-1}(i)$ are directed (by Definition 2.1.1.) and monochromatic (by Definition 2.1.3.).
@@
If $G$ is subgraph of a minimal Cayley graph, then by Observation 2.1 $G$ is the underlying graph of $D$ a subdigraph of a minimal Cayley digraph. Now, we can apply Lemma 2.3 to $D$. By Observation 2.2.2 we get:
@@ 3 An algorithm
3cc: the color currently being extended.
4f_validf\_valid: True if $f^{-1}(c)$ satisfies Definition 2.2, False otherwise.
@@
6if all edges are colored and f_validf\_valid then
@@ 25 $f(e)\leftarrow c$
For the following assume we are given $G=(V,E)$ together with a no-lonely-color coloring $f:E\to S$. We devise an algorithm to construct an orientation $D$ of $G$ color by color to obtain a candidate orientation. It is based on the following easy observations. For this, we note that if $H$ is connected and non-trivial, i.e., with more than one vertex of maximum degree $\leq 2$. Then $H$ has exactly two orientations satisfying Definition 2.1.1, denoted by $H^{+}$ and $H^{-}$. By Observation 2.2 in each color, one of these orientations can be arbitrarily fixed. Further, note that by Observation 2.4, only if some color $f^{-1}(i)$ is a matching, then its edges can potentially be completely bi-oriented. We initialize $D=G$ unoriented and $c=1$.
1Input: An integer $c$ indicating the next color class to orient.
@@ 12      Bi-orient all edges in $f^{-1}(c)$ in $D$
1Input:
2D=(V,A),fD=(V,A),f: A colored directed graph, where $f$ assigns a color to each arc.
@@ 8    $C=(a^{\sigma(1)}_{1},\ldots,a^{\sigma(\ell)}_{\ell})$
9   where each $a_{i}$ is an arc and $\sigma(i)\in\{+1,-1\}$ indicates direction.
10  1exLet $\mathbf{c}=(f(a_{1}),\ldots,f(a_{\ell}))$ be the color sequence.
11   Let $\boldsymbol{\sigma}=(\sigma(1),\ldots,\sigma(\ell))$ be the direction sequence.
12  1exfor every vertex $v$ in $V$ do
@@
4 Two cubic graphs of girth 5
There are two cubic graphs of girth 5. Our implementation [24] applied to the triplex graph, see the left of Figure 1, shows that it is not the subgraph of a minimal Cayely graph. Indeed, Algorithm 1 produces all 216 faithful no-lonely color colorings associated to this graph, in 11 seconds. After reducing modulo automorphism, we are only left with 15 colorings. Algorithm 2 and Algorithm 3 finish without finding any orientation to any of these colorings in about 45 seconds.
@@ 4 Two cubic graphs of girth 5
The second cubic graphs of girth 5 is the twinplex graph depicted on the right of Figure 1. In this case, our procedure finds an orientation and a coloring that passes our algorithm, as depicted in the figure. Hence, we cannot conclude whether the graph is a subgraph of a minimal Cayley graph or not.
@@ 5 Generalized Petersen Graphs
Let $k,n$ be integers such that $0<k<\frac{n}{2}$. The generalized Petersen graph $G(n,k)$ is the cubic graph on vertex set $V=V_{I}\cup V_{O}$, where $V_{I}=\{v_{0},\ldots,v_{n-1}\}$ is the set of inner vertices and $V_{O}=\{u_{0},\ldots,u_{n-1}\}$ the set of outer vertices. The edge set is partitioned into three parts (all subscripts are considered modulo $n$): the edges $E_{O}(n,k)=\{u_{i}u_{i+1}\,|\,0\leq i\leq n-1\}$ inducing a cycle of length $n$; the edges $E_{I}(n,k)=\{v_{i}v_{i+k}\,|\,0\leq i\leq n-1\}$, inducing $\gcd(n,k)$ cycles of length $\frac{n}{\gcd(n,k)}$; and the edges $E_{S}(n,k)=\{u_{i}v_{i}\,|\,0\leq i\leq n-1\}$ forming a perfect matching of $G(n,k)$. Generalized Petersen graphs were introduced by Coxeter [6] and named by Watkins [27]. Many algebraic properties of $G(n,k)$ depend on the particular $k,n$, e.g., isomorphisms [26], automorphism groups, edge-and vertex-transitivity [9], being a core [11], and being a Cayley graph [20, 18].
@@
$a,b\mid a^{r}=b^{n}=e,aba^{-1}=b^{k}>.$
We show that $G(n,k)$ is isomorphic to an induced subgraph of $\textrm{Cay}(C_{n}\rtimes C_{r},\{a,b\})$. Indeed, for $0\leq i\leq n-1$ map every outer vertex $u_{i}$ to $b^{i}$ and every inner vertex $v_{i}$ to $b^{i}a$. See the left side of Figure 2 for an example. Clearly, by minimality of the generating set, the outer vertices induce a cycle of length $n$ in $\textrm{Cay}(C_{n}\rtimes C_{r},\{a,b\})$. Further, there is an edge between $u_{i}$ and $v_{j}$ if and only if $b^{-i}b^{j}a\in\{a,b\}$ or $a^{-1}b^{-j}b^{i}\in\{a,b\}$. This is the case if and only if $i=j$. Now, consider inner vertices $v^{i}v^{j}$ these have an edge if $a^{-1}b^{-i}b^{j}a\in\{a,b\}$ or $a^{-1}b^{-j}b^{i}a\in\{a,b\}$. Using $aba^{-1}=b^{k}$ one gets that this is the case if and only if $i-j=k\mod n$, which finishes the proof. ∎